OpenStudy (shelby1290):

Solve by elimination: 4a/3 - b/4 =6 5a + b = 13

OpenStudy (anonymous):

this Is me

OpenStudy (templeguy):

a=3b16+92

OpenStudy (templeguy):

a=3b/16+92

OpenStudy (templeguy):

number 2 is a=−b5+135

OpenStudy (shelby1290):

@templeguy how did you get a=3b/16+92

OpenStudy (templeguy):

that's number 2 a=-b/5+13/5

OpenStudy (shelby1290):

@Luigi0210 can you please help me with this?

OpenStudy (luigi0210):

Have you tried getting rid of b?

OpenStudy (shelby1290):

No, because I don't understand how to.

OpenStudy (luigi0210):

Okay, so we have \(\Large \frac{4}{3}a - \frac{b}{4}= 6\) \(\Large 5a+b =13\) ---------------------------- So to eliminate b, we need to get of that fraction, so we multiply by 4: \(\Large \frac{16}{3}-b =16 \) \(\Large 5a+b=13 \) --------------------------- Now we can use elimination :P

OpenStudy (shelby1290):

-b=16- 16/3 b=13-5a

OpenStudy (shelby1290):

@Luigi0210 right? ^

OpenStudy (luigi0210):

Whoops, I forgot my a there.. but close enough. When you cancel the b's, you'll end up with: \(\large \frac{16}{3} a=16 \) \(\large 5a=13 \) ------------------------------------------ Now just add them: \(\Large \frac{16}{3}a +\frac{15}{3} a =13+16\) \(\Large \frac{31}{3}a =29 \) ------------------------------------------- This is looking weird but the answer will for a will be.. : \(\Large a=29*\frac{3}{31} \) \(\Large a=\frac{87}{31} \) Hmm, well I guess that's it for a >.<

OpenStudy (shelby1290):

Isn't it supposed to be 16/3b not a ?

OpenStudy (luigi0210):

We're solving for a tho. For elimination, you "eliminate" one of the variables, which was b, so we could solve for the other

OpenStudy (shelby1290):

ohh okay

OpenStudy (shelby1290):

now is the next step to sub in 87,31 into the first equation?

OpenStudy (shelby1290):

a=87/31 **

OpenStudy (luigi0210):

Yup, and that would give you the other variable.

OpenStudy (luigi0210):

Hold on tho, I messed up on the calculations somewhere :/

OpenStudy (shelby1290):

kk..

OpenStudy (luigi0210):

You at least understand the basic steps for solving using elimination right?

OpenStudy (shelby1290):

yea i understand it a little bit

OpenStudy (shelby1290):

would i have to bring 4(87/31)/3 to the other side to make it negative? : -b/4= 6 -4(87/31)/3

OpenStudy (shelby1290):

...to isolate b

OpenStudy (shelby1290):

@Luigi0210 what am i supposed to do next

OpenStudy (luigi0210):

Okay, sorry about that wait.. so attached I have the solution that I came up with. I tried showing the steps. And also, Wolfram confirms it's right :) http://www.wolframalpha.com/input/?i=solve+4a%2F3+-+b%2F4+%3D6+and+5a+%2B+b+%3D+13

OpenStudy (luigi0210):

That should be 152, not 53, sorry

OpenStudy (shelby1290):

Oh cool. So, the poi is (111/31, -152/31)

OpenStudy (shelby1290):

Thanks a millionnnnnnnnnnnn

OpenStudy (shelby1290):

I'm having so much trouble with math lately lol

OpenStudy (luigi0210):

No problem, that's why I'm here to try and help if I can :P

OpenStudy (shelby1290):

Well, thank you for being so caring :)

OpenStudy (luigi0210):

Yup, good luck :D